OpenStudy (anonymous):

a photon of wavelength o.60817 nm strikes a free electron that is initially at rest , the photon is scattered straight backward , what is the speed of the recoil electron after the collision.

OpenStudy (souvik):

 use \[E^2=(mc)^2=(pc)^2+(m _{0}c^2)^2\] formula & conservation of linear momentum and conservation of energy ... eliminate \( \lambda\)' to find the velocity of electron...